### v1

# Trojan Cleansing with Neural Collapse

Abstract
Trojan attacks are sophisticated training-time attacks on neural networks that embed backdoor triggers which force the network to produce a specific output on any input which includes the trigger. With the increasing relevance of deep networks which are too large to train with personal resources and which are trained on data too large to thoroughly audit, these training-time attacks pose a significant risk. In this work, we connect trojan attacks to Neural Collapse, a phenomenon wherein the final feature representations of over-parameterized neural networks converge to a simple geometric structure. We provide experimental evidence that trojan attacks disrupt this convergence for a variety of datasets and architectures. We then use this disruption to design a lightweight, broadly generalizable mechanism for cleansing trojan attacks from a wide variety of different network architectures and experimentally demonstrate its efficacy.
1 Introduction
Over the past decade, deep neural networks have achieved state-of-the-art performance in a vast array of sensitive machine learning (ML) tasks such as autonomous driving, medical diagnostics, and financial portfolio management. And with the advent transformer architectures and large language models, these ML technologies are being incorporated in an ever increasing number of applications and products. The unprecedented speed of this adoption, along with prior research establishing the vulnerability of these methods to a wide variety of attacks, has made the study of security and safety of our ML models an important area of research. This paper considers an important class of training-time attacks on these models known as trojan attacks. In particular, our work attempts to 1) draw upon new results in deep learning theory to understand how trojan attacks alter the structure of neural networks, 2) experimentally validate these conclusions, and 3) use these insights to design and test a lightweight, broadly applicable method for mitigating trojan attacks.
Trojan attacks, or backdoor attacks, aim to maliciously taint the neural network training process so that the trained model so that the presence of the trojan trigger on any input will force the network to produce a target classification [10]. This type of attack is generally accomplished by means of data poisoning, where the adversary adds their chosen trigger to a small portion of a training dataset and labels the poisoned samples with their selected target class. Models trained on this dataset then learn both the correct mappings for clean data as well as the malicious association for inputs containing the trigger. Figure 1 illustrates an example of clean data point (a) along a poisoned one (b). While the yellow trigger in this example is quite visible, as shown by [24], poisoning may take on a subtle nature that is undetectable by human auditors so that that malicious behavior can be hidden until test time or model deployment. This, alongside the incredible amounts of data scraped for use in training very large models, makes trojans a tangible threat even for users training their own models.
Figure 1: An example of trojan attack on a traffic sign classifier. (a) model works normally on clean data without triggers, (b) attackers manipulate the prediction by adding a specific trigger. Image credit: [26].
In this paper, for the first time, we make a concrete connection between training with poisoned data and a phenomenon recently studied in the deep learning theory literature, called Neural Collapse (NC) [22]. Neural Collapse describes the widely observed phenomenon wherein, over the course of training a neural network, highly symmetric structures emerge in both the network’s weights and the feature representations of training samples. In contrast, trojan attacks are inherently asymmetric: they require any input be classified to a single target class given only the presence of a small perturbation. This tension motivates our work here.
1.1 Contributions
While there is a large body of work on trojan detection and mitigation, there has been limited investigation into the mechanism by which trojan triggers act within neural networks with the exception of the topological analysis in [33] observing that the trojaned networks tend to have “skip connections” identified by abnormally strong correlation between neurons in early and late layers in the network. Our analysis is similar in nature as we try to explain and address the problem through an analytic lense.
We pursue this question instead through the inherent asymmetry introduced by training on poisoned datasets and the natural impact this has on the Neural Collapse phenomenon. We discuss NC in depth in Section 2.3, but, briefly, NC is a behavior of over-parameterized neural networks that are trained for an extended period of time after reaching perfect training accuracy. This behavior, associated with generalization [23] and robustness [22], ensures that, when trained on clean datasets, the final feature representations of the training data and the weight matrix of the final layer of the network both converge to a simple, highly symmetric geometric structure. This symmetric structure, however, comes directly in contradiction to the asymmetry introduced by data poisoning. In this work, we present extensive experimental evidence that the process of embedding a trojan in a neural network significantly weakens the NC phenomenon in measurable ways.
We leverage these insights to initiate a new class of trojan cleansing methods: algorithms which take a potentially trojaned network and remove the trojan trigger while maintaining the network’s performance on clean data. Many existing cleansing methods, as detailed in Section 2.2, rely either on some form of model compression, which may harm the model’s generalization capabilities, or on reconstructing the trojan trigger, which relies on knowledge of the class of the triggers the adversary may have chosen from. We are able to instead make a small, carefully targeted adjustment to the weights of the network without any knowledge of the possible trojan triggers.
We conduct extensive evaluations of our cleansing method, comparing with many other standard cleansing algorithms over different network architectures and types of trojan attacks. Our experiments show that our cleansed network effectively maintains the accuracy of the original network on unseen, un-triggered test data, but is no longer susceptible to the trojan trigger. We achieve comparable performance to other cleansing algorithms against standard data poisoning attacks on ResNet architectures and state-of-the-art performance against more sophisticated trojan attacks and large transformer architectures. Additional experiments show our algorithm is particularly robust to data imbalance and corruption. Moreover, our algorithm is very lightweight and easy to implement for any standard classification network architecture. This stands in contrast to many other common cleansing techniques and is particularly important since the trojan threat model assumes that the user must outsource their model training and so likely has limited ML resources and expertise.
2 Related Work
Figure 2: Visualization of NC. The figure illustrates NC as training progresses from left to right. Green spheres represent the vertices of the standard ETF, red spheres and connecting lines represent the linear classifiers, blue spheres and lines represent class means, and small blue spheres represent last-layer features. Each class is distinguished by different color shade. As training progresses, the last-layer features converge onto their respective class means, the class means align with the vertices of the simplex ETF, and the linear classifiers approach (the dual of) their corresponding class means, figure due to [22]
2.1 Trojan Attacks
The baseline data poisoning trojan attack was explored in [5] and [10]. The TrojanNN [17] attack, on a pre-trained model, carefully designs a trigger via network analysis and then reinforces it with a data poisoning step. Wanet [21] introduces warping-based triggers, while ReFool [18] introduces reflection triggers, both offering stealthier triggers than traditional trojan attacks. Instead of using the manually defined trigger, Lira [8] jointly learns a subtle trigger injection function while poisoning the model.
Our work in this paper focuses on the image domain, but trojans have also been proven effective in other domains: [6] and [7] implement trojan attacks on language data and [30] poisons audio data to trojan a model for speech recognition. Extending our analysis and cleansing mechanism to these domains is an interesting direction for future work.
2.2 Model Cleansing
Many approaches to model cleansing task use model compression techniques to remove the portion of the network responsible for the action of the trigger. Li et al. [15] and Xia et al. [29] use forms of distillation to preserve only the model’s performance on clean data while Wang et al. [27] identifies and prunes neurons likely associated with the trojan trigger. Other approaches rely on first reconstructing potential triggers: [4] learns a generative model to create likely triggers and then corrects the network via adversarial training. Zeng et al. [31] proposes a minimax formulation and computationally efficient optimization scheme for trigger reconstruction. Other approaches include [13], which leverages sparse reconstruction of inputs and feature representations to remove potential triggers, and [35], which applies adversarial training to a new layer inserted into the network. [34] is a detection and mitigation method that shrinks the norms of backdoor-related neurons by using sharpness-aware minimization with fine-tuning. Wu et al. [28] introduced BackdoorBench, a platform which implements many of these algorithms.
Related, but distinct work focuses only on detecting whether or not a model may be trojaned. A small selection of such works include those focusd on data analysis, as in Baracaldo et al. [2], Liu et al. [16], Shen et al. [25] and those which directly analyze the model, such as Chen et al. [3], Ma et al. [19], Fields et al. [9].
2.3 Neural Collapse
Neural Collapse is a phenomenon, first reported by Papyan et al. [22] and thoroughly reviewed by [14], which has been observed across a wide variety datasets and model architectures. It describes a process observed in over-parameterized classification deep neural networks which are trained well past achieving zero training error, into a period referred to as the terminal phase of training (TPT). As depicted in Figure 2, as the TPT progresses, the final feature representations of the training data and the weight matrix of the network’s last layer gradually converge to a highly uniform, symmetric geometric structure known as a simplex equiangular tight frame (simplex ETF).
While NC is associated with models trained well past perfect training accuracy, this practice is known to have practical benefits as in the double descent phenomenon [20]. Further experiments in [22] demonstrate in particular that the TPT offers several practical advantages including improved generalization performance, enhanced robustness, and increased interpretability. Various theoretical models, based on locally elastic dynamics [32] and unconstrained features [36, 35] have begun examining how NC arises and its connections to these other neural network characteristics.
It is important to note that Neural Collapse is not a binary phenomenon that either does or does not occur. In practice the characteristics of collapse gradually arise to varying extents over the course of the TPT. Consequently, our discussion of NC will focus on its degree, quantified by four metrics which converge to 0 as the various qualitative characteristics of NC are attained [22]. We will describe these metrics in detail in Section 3.2 after introducing notation.
3 Methodology
3.1 Problem Setting
We consider a deep learning classification problem with $K$ classes, $d-$dimensional training samples $\bm{X}_{i}=\left\{x_{1},\ldots,x_{n_{i}}\right\}$ for each class $i$, and neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{K}$. Our analysis will focus on the final feature representation of the network, so we will decompose the network as $f(\bm{x})=\bm{W}g(\bm{x})+\bm{b}$, with $\bm{W}\in\mathbb{R}^{K\times m}$ and $\bm{b}\in\mathbb{R}^{m}$ giving the final fully-connected layer’s weights and biases, and $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ calculating the network’s final $m-$dimensional feature representation.
The trojan attacks we evaluate in our experiments are carried out via data poisoning. To implement such an attack, we first choose a trigger function $\kappa:\mathbb{R}^{d}\to\mathbb{R}^{d}$. In our experiments this function will be the overlay of a small, solid patch on the base image, as shown in Figure 1. We then choose a poisoning proportion $\delta\in[0,1)$ and a target class $k_{T}\in[K]$. For each $\bm{X}_{k}$ with $k\neq k_{T}$ and each sample $\bm{x}\in\bm{X}_{k}$, with probability $\delta$ we delete $\bm{x}$ from $\bm{X}_{k}$ and add $\kappa(\bm{x})$ to $\bm{X}_{k_{T}}$. We say a model is trojaned if $\delta>0$, and a model is benign if $\delta=0$.
The Neural Collapse phenomenon is observed primarily in two sets of values. The first are the rows of the final layer weights associated with each class, given by $\textbf{W}{\eqqcolon}[\textbf{w}_{1},\ldots,\textbf{w}_{K}]^{T}$, where each row $\textbf{w}_{k}\in\mathbb{R}^{m}$ gives the weights of the final classifier layer associated with class $k$. The second are the centered class means of the final feature representations of the training samples. Define the feature means and the global feature mean:
$\bm{\mu}_{k}\coloneqq\frac{1}{|\bm{X}_{k}|}\sum_{\bm{x}\in\bm{X}_{k}}g(\bm{x}),\quad\bm{\mu}_{G}\coloneqq\frac{1}{K}\sum^{K}_{k=1}\bm{\mu}_{k}.$
Then we can construct the matrix of centered feature means, $\textbf{M}{\coloneqq}[\bm{\mu}_{1}-\bm{\mu}_{G},\ldots,\bm{\mu}_{K}-\bm{\mu}_{G}]^{T}{\eqqcolon}[\textbf{m}_{1},\ldots,\textbf{m}_{K}]^{T}$, and quantify the degree of Neural Collapse in terms of a variety of characteristics of the M and W matrices.
3.2 Neural Collapse Metrics
The observed phenomenon of Neural Collapse can be more precisely described by a set of four characteristics laid out by [22] and quantified by metrics defined in [11]. We will use these quantities to track the degree of NC occurring over training for both benign and trojaned models, and we will show that embedding a trojan trigger in a model significantly slows and weakens NC along multiple dimensions.
NC1: Variability collapse The final feature representation of every training sample in every class converges to the mean of the feature representations of the samples from their respective classes. Equivalently, the variability of feature representations across any given class goes to 0. This can be quantified by defining the within-class and between-class covariances
$\Sigma_{W}$ $\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|\bm{X}_{k}|}\sum_{\bm{x}\in\bm{X}_{k}}\bigg{(}(g(\bm{x})-\bm{\mu}_{k})(g(\bm{x})-\bm{\mu}_{k})^{T}\bigg{)}$
$\Sigma_{B}$ $\coloneqq\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{k}\bm{m}_{k}^{T}.$
We can then quantify the degree of variability collapse by the relationship between the spread of features within a class relative to the spread between the feature means of different classes:
[…]
NC2: Convergence to simplex ETF The rows of both $\bm{M}$ and $\bm{W}$ converge to a simplex ETF. In particular, this means that the norms of each of the rows equalize and that they become maximally pairwise separated from each other as measured by cosine similarity. Note that the maximum pairwise cosine similarity among a set of $K$ vectors is $-(K-1)^{-1}$. These two characteristics can be quantified, respectively, by the two statistics
[…]
We note that this defines the NC2 metric for both the centered class means, $\bm{M}$, and the classifier weights, $\bm{W}$. Our experiments will show that trojan attacks affect both the norm and angular properties of the features and classifier weights in different ways.
NC3: Convergence to self-duality The rows of the final layer weight matrix W converge to a simplex ETF dual to that of the feature representations. Given NC2, this means that $\bm{W}$ and $\bm{M}$ converge, up to a rescaling by their respective Frobenius norms.
[…]
NC4: Nearest feature mean classification The inference time operation of the neural network reduces to a nearest neighbor decision rule in the final feature space. That is, a new test point is classified to the class with the feature mean nearest to its own feature representation. A metric for this is then the misclassification rate for this decision rule.
[…]
$\operatorname*{arg\,min}_{c\in[K]}\left\lVert g(\bm{x})-\bm{\mu}_{c}\right\rVert_{2}\bigg{)}$
Extensive experiments conducted by [22] and [11] demonstrate that NC occurs across a wide variety of different architectures and datasets, with all of the the NC metrics consistently converging towards zero as models progress through the terminal phase of training. However, our observations indicate that application of a trojan trigger weakens this phenomenon. Specifically, under the same model settings and training processes, the NC1-4 metrics for trojaned models are frequently slower to converge than the benign models and often converge to a value larger those of analogous benign models.
(a) Equinorm metric [2] for the feature class means and the final classifier layer weight matrix.
(b) Equiangular metric [2] for the feature class means and the final classifier layer weight matrix.
Figure 3: Convergence to ETF Comparison of the convergence of trojaned and benign models on the NC2 metrics for CIFAR-10 on ResNet18.
3.3 Proposed Trojan Cleansing Method (ETF-FT)
Our experimental results in Section 4.1 strongly suggest that trojan attacks are incompatible with Neural Collapse. Unfortunately, it is difficult to use it to detect trojans–since the scale and rate of convergence of the Neural Collapse mechanisms varies strongly with the network architecture and dataset. Setting an expected degree of collapse associated with a trojan attack would be highly specialized to specific learning problems. Instead, we leverage our insights to design a method for cleansing a network of any potential trojan triggers, which we name ETF-FT.
ETF-FT aims to take a fully trained, possibly trojaned neural network and remove trojan triggers, if any exist, while maintaining performance on clean test data. As specified in Algorithm 1, we proceed by first over-writing the weights of the final fully-connected layer to a randomly generated simplex ETF, freezing those weights, and then fine-tuning the remaining model parameters on a small subset of clean data. Notably, this process works robustly with limited data, and the fine-tuning data does not necessarily need to come from the original training set.
Algorithm 1 Trojan Cleansing with Neural Collapse (ETF-FT)
Input: Trojaned model $h_{\text{troj}}$, clean dataset $\mathbf{X}_{\text{clean}}$
Step 1: Replace weight matrix with $\textbf{W}_{\text{ETF}}$
Construct $\textbf{W}_{\text{ETF}}$ according to Algorithm 2
Replace the final layer weight matrix with $\textbf{W}_{\text{ETF}}$
Step 2: Fine-tune the model
Freeze the final layer weight matrix
Fine-tune on the clean dataset $\mathbf{X}_{\text{clean}}$
Output: Cleansed model $h_{\text{cleansed}}$
Algorithm 2 Constructing the $\mathbf{W}_{\text{ETF}}$ Matrix
Input: Number of classes $K$, random orthogonal matrix $\mathbf{P}\in\mathbb{R}^{m\times K}$
Initialize: $\mathbf{I}\leftarrow\mathbf{I}_{K}$triangleright Identity matrix of size $K\times K$
Initialize: $\mathbf{1}\leftarrow\mathbf{1}_{K\times K}$triangleright Matrix of ones of size $K\times K$
Compute $\mathbf{W}_{\text{ETF}}\leftarrow\sqrt{\frac{K}{K-1}}\times\mathbf{P}\times\left(\mathbf{I}-\frac{1}{K}\mathbf{1}\right)^{T}$
Output: Matrix $\mathbf{W}_{\text{ETF}}$
ETF-FT is motivated by our results which show that trojans work, in part, by deforming the symmetric structure NC describes in the weights of the final layer. By fixing those weights to be a random, but perfectly symmetric ETF we accomplish two things: first, we prevent the trojan deformation of the final layer and second, we reset the weights in a way known to be effective for classification. This addresses a challenge present in many trojan cleansing techniques: the more you perturb the poisoned model the less effective the trigger will be, but the harder is it to recover accuracy on clean data.
3.4 Experiment Details for Cleansing Methods Comparison
We compare our method with several state-of-the-art cleansing algorithms. Our evaluation considers two types of backdoor attacks: BadNets-A2O [10] and TrojanNN [17], and two model architectures: PreAct-ResNet18 [12] and ViT-B/16 [1].
We focus on two key metrics to evaluate the performance of each cleansing method: accuracy on clean data (ACC) and attack success sate (ASR). Accuracy reflects the model’s performance on clean data after cleansing, ideally showing no reduction compared to the original model. ASR, on the other hand, measures the percentage of poisoned samples classified into the attack target class. An effective cleansing method should reduce ASR to nearly zero.
We utilize the BackdoorBench repository [28] implementations of all attacks and cleansing methods besides our own. For experiments on PreAct-ResNet18, we use the original hyperparameters provided in the repository. By default, the repository uses 5% of the training data as available clean data for fine-tuning, though we also run experiments with 1% of training data and with imbalanced and corrupted data, as in real world situations the end user may not have access to (much of) the original training data.
We also evaluate the cleansing algorithm’s performance on a vision transformer, due to the popularity and efficacy of transformer architectures. However, many cleansing algorithms are not designed with transformers in mind. Where necessary, we made design choices, as in NPD, which inserts a new layer into the network, we chose to insert a linear layer before the final layer of the transformer encoders. We also did mild training hyperparameter tuning over batch size, learning rate, scheduler and decay rate for each algorithm including our own. Despite this, many algorithms show poor performance in this setting, relative to their results on ResNets. We want to emphasize that many of these algorithms are likely capable of better performance on transformers given careful adaptation. However, in the trojan attack scenario, the end user has outsourced network training to an adversary and is therefore unlikely to be a sophisticated deep learning engineer. As such, it is a significant strength of our algorithm that it can be applied to any classifier architecture without modification.
4 Results
4.1 Neural Collapse Disruption
CIFAR-100 GTSRB
Benign 1% 5% Benign 5% 10% 20%
NC1 0.8917 0.9165 0.8830 0.1567 1.3484 2.2541 5.0615
[…]
NC3 0.0174 0.0183 0.0254 0.1419 0.1747 0.2115 0.2376
NC4 0.0000 0.0000 0.0001 0.0000 0.0174 0.0691 0.1680
Table 1: Comparison of all NC metrics between trojaned and benign models for CIFAR-100 and GTSRB. $\bm{M}$ refers to the matrix of class means of the final-layer feature representations; $\bm{W}$ denotes the weight matrix of the final fully-connected layer. The percentage indicates the poison rate for the training dataset of the trojaned models.
We first examine the behavior of the NC metrics described in Section 3.2 for both trojaned and benign models over the course of training on the CIFAR-10, CIFAR-100, and GTSRB datasets. Our results generally show that the application of a trojan trigger causes the NC metrics to converge more slowly and to converge to a larger value than their benign counterparts. There is variation in this behavior across different datasets and different metrics, but in every case the trojaned model shows weaker collapse in at least some metrics.
Figure 3 tracks NC2, defined in Equation 2, which characterizes the convergence of both the final classifier weights and the training feature means to a simplex ETF. The red dashed line in the figures marks the point where the training error reaches zero and so the regime in which Neural Collapse is generally observed. For the equinorm property, shown in Figure 3(a), the benign model shows the most uniformity in norm across both feature means and classifier weights, and as the percentage of trojaned data increases, the equinorm property weakens. This trend is particularly profound in the classifier weights. We note that, for both class means and classifiers, the source of this deviation is actually that the norm of the target class is consistently smaller than those of other classes. For the equiangular comparison, a similar, though smaller trend is observed. In Figure 3(b), the benign model reaches the smallest equiangular metric for both feature class means and classifiers. This further underscores the weakening of Neural Collapse due to trojaning.
NC3, defined in Equation 3, is shown in Figure 4. Here, again, the metric converges to a smaller value for the benign model than either trojaned model and the disruption to Neural Collapse appears to increase with the proportion of training data poisoned with the trojan trigger.
Figure 4: Self-duality Comparison of the convergence of trojaned and benign models on the NC3 metric [3] for CIFAR-10 on ResNet18.
But this pattern is not observed in every collapse metric for every dataset. On CIFAR-10 we did not observe a significant difference between the benign and trojaned models for NC1, characterizing the variability across feature representations, or NC4, characterizing how close the network’s decision rule is to a nearest neighbor rule in the final feature space, as they all go down to zero. However, collapse is noticeably weakened in these metrics in our experiments on CIFAR-100 and GTSRB, as shown in Figures 6 and 6 and Table 1.
Figure 5: Variability Collapse Comparison of the convergence of trojaned and benign models on the NC1 metric [1] for GTSRB on ResNet18.
Figure 6: Nearest feature classification Comparison of the convergence of trojaned and benign models on the NC4 metric [4] for GTSRB on ResNet18.
PreAct-ResNet18[12] ViT-B/16[1]
5% Clean data 1% Clean data 5% 1%
Method ACC ASR ACC ASR ACC ASR ACC ASR
Original 88.95 95.06 88.95 95.06 95.43 87.06 95.43 87.06
FT $91.63\pm 0.18$ $10.03\pm 7.23$ $90.20\pm 0.22$ $22.04\pm 25.38$ 95.43 87.06 95.74 36.87
NCleanse[27] $91.57\pm 0.1$ $\textbf{0.98}\pm 0.12$ $89.90\pm 0.34$ $\textbf{1.12}\pm 0.18$ 96.08 0.28 95.61 2.70
NAD[15] $91.30\pm 0.18$ $2.20\pm 0.43$ $88.88\pm 0.28$ $10.19\pm 12.98$ 94.25 2.12 95.24 29.42
NPD[35] $85.71\pm 3.59$ $3.39\pm 2.47$ $83.26\pm 3.59$ $5.26\pm 3.82$ 91.38 50.35 91.92 85.17
FT-SAM[34] $\textbf{92.23}\pm 0.12$ $2.37\pm 0.81$ $\textbf{90.33}\pm 0.15$ $8.12\pm 11.86$ 96.38 1.93 94.29 6.16
I-BAU[31] $88.78\pm 1.53$ $5.53\pm 6.81$ $86.16\pm 0.91$ $32.79\pm 18.81$ 25.96 10.87 12.91 68.68
ETF-FT(ours) $91.77\pm 0.16$ $2.85\pm 0.68$ $90.24\pm 0.28$ $4.22\pm 2.05$ 91.99 1.15 87.55 1.60
Table 2: Comparison of cleansing algorithms on trojaned models trained on 10% poisoned data with the BadNet attack. Results on ResNet18 are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
PreAct-ResNet18[12] ViT-B/16[1]
5% Clean data 1% Clean data 5% 1%
Method ACC ASR ACC ASR ACC ASR ACC ASR
Original 91.17 99.90 91.17 99.90 96.67 100.00 96.67 100.00
FT $91.92\pm 0.15$ $3.12\pm 3.14$ $90.60\pm 0.37$ $3.75\pm 6.41$ 95.94 100.00 95.98 100.00
NCleanse[27] $91.49\pm 0.45$ $41.07\pm 48.04$ $90.50\pm 1.05$ $70.67\pm 44.64$ 96.80 2.33 96.52 40.96
NAD[15] $91.36\pm 0.28$ $1.51\pm 0.37$ $88.39\pm 0.89$ $2.99\pm 1.94$ 89.99 100.00 96.40 100.00
NPD[35] $82.65\pm 2.14$ $16.34\pm 11.34$ $86.87\pm 0.13$ $43.03\pm 1.68$ 96.69 100.00 96.50 100.00
FT-SAM[34] $90.78\pm 0.21$ $8.10\pm 4.62$ $87.52\pm 0.60$ $12.23\pm 19.80$ 96.98 100.00 96.82 100.00
I-BAU[31] $91.83\pm 0.12$ $82.22\pm 1.73$ $\textbf{91.73}\pm 0.25$ $99.91\pm 0.01$ 36.10 9.74 36.70 0.41
ETF-FT(ours) $\textbf{92.39}\pm 0.24$ $\textbf{0.04}\pm 0.05$ $91.01\pm 0.47$ $0\textbf{0.00}\pm 0$ 0.00 95.34 10.76 95.04 10.11
Table 3: Comparison of cleansing algorithms on trojaned models trained with 10% poisoned data with the TrojanNN attack. Results on ResNet are averaged over 10 random trials with mean±$\pm$±std; result on ViT is a single trial.
So, while not every metric clearly differentiates trojan and benign models on every dataset, it is the case that benign models always show at least the same degree of convergence as any trojaned model on every collapse metric. And, importantly, on every dataset we observe at least one metric where trojaned models show a significantly weaker degree of collapse than their benign counterpart–suggesting strong tension between Neural Collapse and an effective trojan trigger.
4.2 Trojan Cleansing
We compare our cleansing algorithm, laid out in Section 3.3, with several SOTA cleansing methods on trojaned models. Each of these algorithms uses a small amount of clean data to either analyze or fine-tune the target model. 5% of the training dataset is a common allowance for this setting, but we also run evaluations with 1% of the training dataset as the user in a trojan threat model can be expected to have limited data access. We also use vanilla fine-tuning on the provided dataset as a baseline–these results labeled FT in the following tables.
Our evaluation considers two common trojan attacks: BadNet [10] and TrojanNN [17], and two distinct model architectures: PreAct-ResNet18 [12] and ViT-B/16 [1]. All experiments here were carried out with the BackdoorBench repository [28]. For ResNet, we use the original hyperparameters provided in the configuration. For ViT, we perform moderate hyperparameter tuning for each cleansing method to ensure stable running and reasonable performance. As discussed in Section 3.4, these algorithms may perform better on the transformer models given careful design adaptations, but it is a strength of our algorithm that it readily applies to any classifier architecture out of the box.
BadNet Attack Table 2 presents the cleansing results for the BadNet attack. When the cleansing algorithms are allowed to use 5% of the training data to fine-tune, on the ResNet model, ETF-FT achieves competitive ASR while maintaining high accuracy on clean data, relative to the other methods tested. In more restrictive conditions, where only 1% of the training data is available, ETF-FT and Neural Cleanse maintain strong performance, while other methods show significant degradation.
Vision Transformers Despite their state-of-the-art performance and growing popularity, vision transformers have rarely been examined in the context of backdoor cleansing performance. This is in part due to the fact that their distinct architecture makes it difficult to adapt many common cleansing methods. Table 2 shows cleansing results for the BadNet attack applied to the ViT model, ETF-FT achieves very low ASR in both standard and limited data conditions, surpassing other methods except Neural Cleanse in the standard data case. This does come at the cost of a trade-off with in clean accuracy, which we believe can be mitigated, as discussed in Section 5.1.
TrojanNN Attack Alongside the standard BadNet attack, we consider the more sophisticated TrojanNN attack [17] which identifies specific vulnerable neurons in a pre-trained network and reverse engineers a subtle trigger to target them. As shown in Table 3, ETF-FT is particularly well suited to cleansing this style of trojan. On ResNets, ETF-FT completely eliminates the trigger, achieving an ASR of nearly 0 across 10 trials, while maintaining competitive accuracy with the other cleansing methods. On ViT, many methods, implemented out of the box, fail completely to eliminate the trigger, while ETF-FT achieves an ASR of 10% with only a 1.5 point drop in clean accuracy. The I-BAU algorithm is able to almost entirely eliminate the trigger, but in doing so destroyed the model’s accuracy on clean data.
4.3 Robustness
The above results show that ETF-FT performs well with access to only a small amount of training data. But it is likely that the user of a trojaned model may not have access to any data from the original training distribution at all. We simulated distribution shift in two ways: by imbalancing the class distribution and by applying random erasures within image patches. Results for these datasets, applied to the ResNet BadNet poisoned models, are shown in Tables 5 and 4. In both settings, ETF-FT excels at maintaining the model’s clean accuracy along with competitive ASR relative to other methods. Neural Cleanse, which had generally very strong results in this setting with training data, performs significantly better than all other methods on the imbalanced data, but its clean accuracy suffers on the corrupted data. This is due to the fact that its pruning method relies on identifying important neurons based on their activations on the available data–so it is vulnerable to accidentally pruning neurons associated with valid features if those features aren’t represented in the fine-tuning dataset.
Method ACC ASR
Original 88.95 95.06
FT $81.66\pm 0.39$ $1.76\pm 0.48$
NCleanse[27] $81.80\pm 0.38$ $1.72\pm 0.38$
NAD[15] $67.14\pm 1.96$ $3.10\pm 1.68$
NPD[35] $83.10\pm 3.64$ $5.97\pm 4.59$
FT-SAM[34] $85.23\pm 0.77$ $\textbf{1.21}\pm 0.48$
I-BAU[31] $86.65\pm 1.38$ $28.26\pm 17.91$
ETF-FT(ours) $\textbf{87.99}\pm 0.49$ $3.37\pm 2.80$
Table 4: Data Corruption: Comparison of cleansing algorithms with 1% corrupted (under random erasure) clean training data on BadNet trojaned ResNet models. Results are averaged over 5 random trials with mean±$\pm$±std.
Method ACC ASR
Original 88.95 95.06
FT $86.01\pm 1.59$ $22.95\pm 19.81$
NCleanse[27] $\textbf{88.43}\pm 0.29$ $\textbf{1.19}\pm 0.27$
NAD[15] $83.41\pm 1.87$ $3.17\pm 0.72$
NPD[35] $83.10\pm 3.64$ $5.97\pm 4.59$
FT-SAM[34] $87.20\pm 0.54$ $7.17\pm 3.69$
I-BAU[31] $85.97\pm 1.43$ $49.55\pm 12.15$
ETF-FT(ours) $88.16\pm 0.56$ $7.18\pm 2.06$
Table 5: Data Imbalance: Comparison of cleansing algorithms with 1% imbalanced clean training data on BadNet trojaned ResNet models(%). Results are averaged over 5 random trials with mean±$\pm$±std.
5 Conclusion
In this paper, we explore how the asymmetry induced by a trojan trigger in a neural network disrupts the phenomenon of Neural Collapse. We provide experimental evidence showing that trojan attacks significantly weaken Neural Collapse based on standard metrics. Leveraging these insights, we propose a novel method for trojan cleansing: by fixing the network’s final layer to a simplex ETF—an ideal structure predicted by Neural Collapse—and fine-tuning on a small amount of clean data. Our results show that this approach effectively removes the trojan trigger while maintaining the network’s accuracy on clean test data without requiring knowledge of the trigger type, the target class, or access to the original training data. Moreover, our method can be applied to any standard architecture with no modification.
5.1 Future Work
This work suggests a variety of interesting future directions. While we’ve shown significant experimental evidence of the connection between trojans and Neural Collapse, it would be valuable to develop a theoretical model of their interaction. Experimentally our results could be extended to different architectures, attack types, and trigger types. In particular, a thorough enough characterization of Neural Collapse in this different settings could allow for the use of our results for detection, not just cleansing, of trojan triggers. And, for the efficacy of our cleansing method, we hope to further explore the choice of the ETF imposed on the network. In this work we used a randomly generated ETF, but, intuitively, an ETF nearer that of the network’s feature means would better maintain clean accuracy, but potentially make it harder to reduce trigger efficacy. We believe careful selection of the ETF there may improve the performance of our cleansing mechanism.